# On the existence of weighted-cscK metrics

### v2

@@ 2.C. Examples
After a normalization, $\log(\mathrm{Vol}^{\lambda})$ has the same expression as Perelman’s $W$-functional [Per02]. When restricted on Fano manifolds, $\log(\mathrm{Vol}^{\lambda})$ is equal to Tian–Zhu’s volume functional [TZ02] up to the difference of a constant.
Some other related works can be found in [He19, Ino21, Liu22, SD22, Jub23, Zha23, Fut24, GJSS24, His24, NN24].
@@ 3. Apriori estimates
In this section, we provide uniform $C^{0}$-estimate, $W^{2,p}$-estimate of $\phi$ and gradient estimate of $F+f_{*}-\log(\mathrm{v}(m_{\phi}))$ for equation (2.20a), (2.20b). The proof of the uniform $C^{0}$-estimate is based on a modification of Guo–Phong’s method [GP22]. The proof of $W^{2,p}$-estimate and gradient estimate follows from Chen–Cheng’s method [CC21b] with a new choice of test functions and integral weights.
@@ 3.A. Uniform $C^{0}$-estimate
Let $\omega_{X}$ be a fixed smooth Kähler form, and let $\omega=\chi+s\omega_{X}$ be a smooth Kähler form, where $\chi$ is a smooth $(1,1)$-form in a nef class, $s>0$. We will first consider the uniform $C^{0}$-estimate of the following PDE system
$\begin{cases}\mathrm{v}(m_{\phi})\;\omega_{\phi}^{n}=c_{\omega,\mathrm{v}}\;e^{\widetilde{F}}\omega_{X}^{n},\\ (\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})\big{(}\widetilde{F}+f_{*}\big{)}=-f+\mathrm{tr}_{\phi}\eta,\end{cases}$ (3.1)
@@
In the modified continuity path, for the $(\frac{1-t}{t})\theta$-psh function $f_{*}$, for arbitrarily large $p$, we have $\int_{X}e^{-pf_{*}}\omega^{n}$ uniformly bounded from above when $t$ is sufficiently close to $1$ (see Proposition 4.8). Furthermore, the function $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is determined by the class $[\theta]$, which is bounded. Therefore the modified continuity path (2.20a), (2.20b) can be reformulated into (3.1) when $t$ is sufficiently close to $1$.
@@
By applying [GP22, Theorem 2] to equation $\omega_{\phi}^{n}=c_{\omega}e^{\widetilde{F}}\omega_{X}^{n}$, for $l>n$, we have $\sup_{X}|\phi|\leq C$, where $C$ depends on $\omega_{X},n,l,\kappa,{\rm Ent}_{l}(e^{\widetilde{F}})$.
Therefore, we just need to modify [GP22, Theorem 3] to use $\int_{X}\widetilde{F}e^{\widetilde{F}}\omega_{X}^{n}$ to bound ${\rm Ent}_{l}(\widetilde{F})$.
@@
where $\Omega=\{x\in X:\widetilde{F}+f_{*}>0\}$ and $0\leq A_{\infty}\leq C(K_{1})$, which is due to the entropy bound and the weak compactness implied by it.
@@
where $\epsilon=(\frac{(n+1)(n+C_{f,\delta})}{n^{2}})^{\frac{n}{n+1}}A_{k}^{\frac{1}{n+1}}$, $\Lambda^{\frac{1}{n+1}}=\frac{2n}{n+1}\epsilon$, where the constant $C_{f,\delta}$ is chosen such that $\left|-\delta f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)\right|\leq C_{f,\delta}$.
@@
which implies
$\delta\cdot\Delta_{\phi}(\widetilde{F}+f_{*})=-\delta\cdot f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)+\delta\cdot\mathrm{tr}_{\phi}\eta-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})\cdot(-\psi_{k}+\Lambda)^{-\frac{1}{n+1}}.$
@@
$\quad+\left(1-\frac{\epsilon n}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}-\frac{1}{10}\right){\rm tr}_{\phi}\omega-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}$
@@
$-\phi+\delta(\widetilde{F}+f_{*})\leq\left(\frac{(n+1)(n+C_{f,\delta})}{\epsilon n^{2}}\right)^{n}A_{k}(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+C^{\prime}=\epsilon(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+C^{\prime}.$
@@
In particular, this implies $\mathrm{Ent}_{l}(\widetilde{F})$ is uniformly bounded. Then by [GPT23, Theorem 1] we have $\sup_{X}|\phi|\leq C$, where the constant $C$ depends on parameters listed in the statement of the theorem. (Alternatively, we can also use the $C^{0}$-estimate in [EGZ09]).
In order to prove the upper and lower bounds of $\widetilde{F}+f_{*}$, we need to use the mean-value type inequality proved in [GP22].
Let $u:=\widetilde{F}+f_{*}-K_{2}\phi$, which satisfies $(\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})u\geq-a$ for some constant $a\geq 0$. Apply [GP22, Lemma 3] to $u$, we can obtain the upper bound of $\widetilde{F}+f_{*}$.
@@ 3.B. $W^{2,p}$-estimate
$\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}=\Delta_{\phi}f_{*}+\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),\mathrm{d}^{c}f_{*}\rangle}{\mathrm{v}}=\Delta_{\phi}f_{*}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}f_{*},$
@@
$\widetilde{f}_{t}=-(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}+(1-\frac{1}{t})\frac{\langle(\mathrm{d}\mathrm{v})(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}$
@@
$\qquad\quad+\Delta F^{\prime}-S(\omega)\Big{)}+2au|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-au\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+u^{-1}|\nabla^{\phi}u|_{\phi}^{2}\ ,$ (3.9)
@@
Take the integral on the above equality with respect to the weighted measure $e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$, together with (3.9), we obtain
$\int_{X}2pu^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$=$ $\int_{X}u^{2p}\Big{(}-\Delta_{\phi}u-\langle\nabla_{\phi}u,2a\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}\Big{)}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ $-2a\int_{X}u^{2p}\langle\nabla_{\phi}u,\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
@@
$-2a\int_{X}u^{2p+1}|\nabla_{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}-\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
@@
$(2p+1-a)\int_{X}u^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
@@
$-2a\int_{X}u^{2p}\langle\nabla_{\phi}u,\nabla_{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ ${a}\int_{X}u^{2p+1}|\nabla_{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+a\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}.$
@@
$\leq$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-2}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\cdot|\nabla_{\phi}u|_{\phi}^{2}(n+\Delta\phi)e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$=$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-1}|\nabla_{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n},$
@@
$\int_{X}(2p+1-a)u^{2p-1}\cdot|\nabla_{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+\frac{ab}{2}\int_{X}u^{2p+1}\mathrm{tr}_{\phi}\omega e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$

### v3

@@ 2.C. Examples
After a normalization, $\log(\mathrm{Vol}^{\lambda})$ has the same expression as Perelman’s $W$-functional [Pe02]. When restricted on Fano manifolds, $\log(\mathrm{Vol}^{\lambda})$ is equal to Tian–Zhu’s volume functional [TZ02] up to the difference of a constant.
Some other related works can be found in [He19, Ino21, Liu22, Dyr22, Jub23, Zha23, Fut24, GJSS24, His24, NN24].
@@ 3. Apriori estimates
In this section, we provide uniform $C^{0}$-estimate, $W^{2,p}$-estimate of $\phi$ and gradient estimate of $F+f_{*}-\log(\mathrm{v}(m_{\phi}))$ for equation (2.21a), (2.21b). The proof of the uniform $C^{0}$-estimate is based on a modification of Guo–Phong’s method [guo2022]. The proof of $W^{2,p}$-estimate and gradient estimate follows from Chen–Cheng’s method [CC21b] with a new choice of test functions and integral weights.
@@ 3.A. Uniform $C^{0}$-estimate
Let $\omega_{X}$ be a fixed smooth Kähler form, and let $\omega=\chi+s\omega_{X}$ be a smooth Kähler form, where $\chi$ is a smooth $(1,1)$-form in a nef class, $s>0$. We simplify the equations (2.21a), (2.21b). The equation (2.21a) can be rewritten as
$\mathrm{v}(m_{\phi})\;\omega_{\phi}^{n}=e^{F+\log\frac{\omega^{n}}{\omega_{X}^{n}}}\omega_{X}^{n}=e^{\widetilde{F}+\log c_{\omega,\mathrm{v}}}\omega_{X}^{n},$
where we denote $\widetilde{F}=F+\log\frac{\omega^{n}}{\omega_{X}^{n}}-\log c_{\omega,\mathrm{v}}$. On the other hand, the equation (2.21b) can be re-written as
$(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha})(F+f_{*})=-f_{t}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})+\mathrm{tr}_{\phi}((1-\frac{1}{t})\theta+\mathrm{Ric}(\omega)).$
Thus, by (2.12), one has
$(\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha})(\widetilde{F}+f_{*})=$ $-f_{t}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})+\mathrm{tr}_{\phi}((1-\frac{1}{t})\theta+\mathrm{Ric}(\omega))$
$+\Delta_{\phi}\log\frac{\omega^{n}}{\omega_{X}^{n}}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}\log\frac{\omega^{n}}{\omega_{X}^{n}}$
$=$ $-f_{t}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega_{X}}(m_{\omega_{X}}^{\xi^{\alpha}})+\mathrm{tr}_{\phi}\Big((1-\frac{1}{t})\theta+\mathrm{Ric}(\omega_{X})\Big).$
We will consider the uniform $C^{0}$-estimate of the following PDE system
$\begin{cases}\mathrm{v}(m_{\phi})\;\omega_{\phi}^{n}=c_{\omega,\mathrm{v}}\;e^{\widetilde{F}}\omega_{X}^{n},\\ (\Delta_{\phi}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha})\big(\widetilde{F}+f_{*}\big)=-f+\mathrm{tr}_{\phi}\eta,\end{cases}$ (3.1)
@@
In the modified continuity path, for the $(\frac{1-t}{t})\theta$-psh function $f_{*}$, for arbitrarily large $p$, we have $\int_{X}e^{-pf_{*}}\omega^{n}$ uniformly bounded from above when $t$ is sufficiently close to $1$ (see Proposition 4.8). Furthermore, the function $\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}(f_{*})$ is bounded by a constant which only depends on the class $\theta]$. Indeed, denote $\theta_{t}:=\frac{1-t}{t}\theta$ and by (2.9),
$-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}(f_{*})=\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}\mathrm{d}^{c}f_{*}(\xi^{\alpha})=\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}(m_{\theta_{t}+{\rm dd^{c}}f_{*}}^{\xi^{\alpha}}-m_{\theta_{t}}^{\xi^{\alpha}}),$
where the right hand side only depends on the moment polytope determined by $\theta]$.
Therefore the modified continuity path (2.21a), (2.21b) can be reformulated into (3.1) when $t$ is sufficiently close to 11.
@@
[…]
@@
where $\Omega=\{x\in X:-\phi+\delta(\widetilde{F}+f_{*})>0\}$ and $0\leq A_{\infty}\leq C(K_{1})$, which is due to the entropy bound assumption, [guo2022, Theorem 1 (ii)] for $p=1$ and Hölder inequality.
@@
where $\epsilon=2^{\frac{n-1}{n+1}}(\frac{(n+1)(n+C_{f,\delta})}{n^{2}})^{\frac{n}{n+1}}A_{k}^{\frac{1}{n+1}}$, $\Lambda^{\frac{1}{n+1}}=\frac{2n}{n+1}\epsilon$, where the constant $C_{f,\delta}$ is chosen such that $\left|-\delta f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)\right|\leq C_{f,\delta}$.
@@
which together with the second equation of (3.1) implies
$\delta\cdot\Delta_{\phi}(\widetilde{F}+f_{*})=-\delta\cdot f+\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\phi)+\delta\cdot\mathrm{tr}_{\phi}\eta-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})\cdot(-\psi_{k}+\Lambda)^{-\frac{1}{n+1}}.$ (3.4)
@@
$\qquad+\left(1-\frac{\epsilon n}{n+1}(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}-\frac{1}{10}\right){\rm tr}_{\phi}\omega-\frac{n\epsilon}{n+1}\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha}(\psi_{k})(-\psi_{k}+\Lambda)^{\frac{-1}{n+1}}$
@@
$-\phi+\delta(\widetilde{F}+f_{*})\leq$ $\left[C^{\prime}+\frac{(n+1)(n+C_{f,\delta})}{\epsilon n^{2}}(-\psi_{k}+\Lambda)^{\frac{1}{n+1}}\right]^{n}A_{k}$
$\leq$ $2^{n-1}\left(\frac{(n+1)(n+C_{f,\delta})}{\epsilon n^{2}}\right)^{n}A_{k}(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+2^{n-1}C^{\prime n}A_{k}$
$=$ $\epsilon(-\psi_{k}+\Lambda)^{\frac{n}{n+1}}+C^{\prime}.$
@@
[…]
In order to prove the upper and lower bounds of $\widetilde{F}+f_{*}$, we need to use the mean-value type inequality proved in [guo2022].
Let $u:=\widetilde{F}+f_{*}-K_{2}\phi$, which satisfies $(\Delta_{\phi}-\frac{\mathrm{v}_{\alpha}}{\mathrm{v}}J\xi^{\alpha})u\geq-a$ for some constant $a\geq 0$. Apply [guo2022, Lemma 3] to $u$, we can obtain the upper bound of $\widetilde{F}+f_{*}$.
@@ 3.B. $W^{2,p}$-estimate
$\frac{\mathrm{tr}_{\mathrm{v},\phi}({\rm dd^{c}}f_{*})}{\mathrm{v}(m_{\phi})}=\Delta_{\phi}f_{*}+\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),\mathrm{d}^{c}f_{*}\rangle}{\mathrm{v}}=\Delta_{\phi}f_{*}-\frac{\mathrm{v}_{,\alpha}(m_{\phi})}{\mathrm{v}(m_{\phi})}J\xi^{\alpha}f_{*},$
@@
$\widetilde{f}_{t}=-(\ell_{\mathrm{ext}}+(1-\frac{1}{t})\underline{\theta})\frac{\mathrm{w}}{\mathrm{v}}+(1-\frac{1}{t})\frac{\langle\mathrm{d}\mathrm{v}(m_{\phi}),m_{\theta}\rangle}{\mathrm{v}}-\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}m_{\omega}^{\xi^{\alpha}}$
@@
Note that by (2.B) and (2.13), the second equation of (3.5) can be rewritten as
$\Delta_{\phi}(F+\log\mathrm{v}(m_{\phi}))=$ $\widetilde{f_{t}}+\mathrm{tr}_{\phi}(\eta_{t})-\Delta_{\phi}f_{*}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}f_{*}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}F+\Delta_{\phi}\log\mathrm{v}(m_{\phi})$
$=$ $\widetilde{f_{t}}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}\Delta_{\omega}(m_{\omega}^{\xi^{\alpha}})+\mathrm{tr}_{\phi}(\eta_{t})-\Delta_{\phi}f_{*}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}f_{*}$
$-2|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}$
$=$ $-f_{t}+\frac{\mathrm{v}_{,\alpha}}{\mathrm{v}}J\xi^{\alpha}f_{*}-\Delta_{\phi}f_{*}+\mathrm{tr}_{\phi}\eta_{t}-2|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}+\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}.$ (3.6)
@@
$\hskip 18.49988pt\qquad+\Delta F^{\prime}-S(\omega)\Big)+2au|\nabla^{\phi}\log\mathrm{v}(m_{\phi})|_{\phi}^{2}-au\frac{\mathrm{v}_{,\alpha\beta}(m_{\phi})}{\mathrm{v}(m_{\phi})}\langle\xi^{\alpha},\xi^{\beta}\rangle_{\phi}+u^{-1}|\nabla^{\phi}u|_{\phi}^{2}\ ,$ (3.11)
@@
Integrate the equality above with respect to the weighted measure $e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$, together with (3.11), we obtain
$\int_{X}2pu^{2p-1}\cdot|\nabla^{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$=$ $\int_{X}u^{2p}\Big(-\Delta_{\phi}u-\langle\nabla^{\phi}u,2a\nabla^{\phi}\log\mathrm{v}\rangle_{\phi}\Big)\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ $-2a\int_{X}u^{2p}\langle\nabla^{\phi}u,\nabla^{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
@@
$-2a\int_{X}u^{2p+1}|\nabla^{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}-\int_{X}u^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
@@
$(2p+1-a)\int_{X}u^{2p-1}\cdot|\nabla^{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
@@
$-2a\int_{X}u^{2p}\langle\nabla^{\phi}u,\nabla^{\phi}\log\mathrm{v}\rangle_{\phi}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$\leq$ ${a}\int_{X}u^{2p+1}|\nabla^{\phi}\log\mathrm{v}|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+a\int_{X}u^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n}.$
@@
$\leq$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-2}e^{-a(F+\log\mathrm{v}+\delta f_{*}+b\phi)}\cdot|\nabla^{\phi}u|_{\phi}^{2}(n+\Delta\phi)e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$
$=$ $\frac{p^{2}}{a-1}\int_{X}u^{2p-1}|\nabla^{\phi}u|_{\phi}^{2}e^{2a\log\mathrm{v}}\omega_{\phi}^{n},$
@@
$\int_{X}(2p+1-a)u^{2p-1}\cdot|\nabla^{\phi}u|_{\phi}^{2}\cdot e^{2a\log\mathrm{v}}\omega_{\phi}^{n}+\frac{ab}{2}\int_{X}u^{2p+1}\mathrm{tr}_{\phi}\omega e^{2a\log\mathrm{v}}\omega_{\phi}^{n}$